The ratio of the sides of a rectangle are 3:5 and it's area is 60 units.Make an equation to solve.
Answer by jim_thompson5910(35256)

3x*5x = 60

15x^2 = 60

x^2 = 60/15

x^2 = 4

x = sqrt(4)

x = 2

-------------------------------------------------------

3x = 3*2 = 6

5x = 5*2 = 10

-------------------------------------------------------

The rectangle has a length of 10 and a width of 6

Notice how 6:10 reduces to 3:5 (divide both parts by 2)